OpenStudy (anonymous):

How many ways are there to choose a committee of 7 people from a group of 10 people?

OpenStudy (anonymous):

Since order of the 7 people does NOT count, you are dealing with a combination. So you want 10 C 7, a combination of 7 from 10 people. So compute 10 C 7.

OpenStudy (anonymous):

how do i compute 10 C 7

OpenStudy (anonymous):

Does your teacher allow you to use a calculator to compute 10 C 7?

OpenStudy (anonymous):

yea but how would i put that in the calcalator ?

OpenStudy (anonymous):

do you know where the combination function, C, is on your calculator? It's probably in your probability menu. so you type 10 C 7, then enter.

OpenStudy (anonymous):

\[_{10}\text{C}_{7} =\frac{ 10! }{ 7!\,(10-7)! }=\frac{ 10! }{ 7!\,3! }\]